# IRS on orbits and the case of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$

### v1

@@ Abstract.
Let $G$ be a connected Lie group and $\text{Sub}_{G}$ be the space of closed subgroups of $G$ equipped with the Chabauty topology. In this article, we investigate the existence of invariant random subgroups of $G$ supported on various orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. We also find the orbit closure for each of these orbits in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
Key words and phrases:
Connected Lie groups, spaces of closed subgroups, Chabauty topology, invariant random subgroups, conjugacy limits
1991 Mathematics Subject Classification:
Primary: $22$E15151515, $28$D15151515; secondary: $37$A15151515, $37$B02020202
@@ 1. Introduction
Let $G$ be a connected (real) Lie group and ${\rm Sub}_{G}$ be the space of (closed) subgroups of $G$ equipped with the Chabauty topology introduced by Claude Chabauty in [4]. The Chabauty topology on SubG is generated by the subsets $\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$, $\{H\in{\rm Sub}_{G}\mid H\cap U\not=\emptyset\}$, where $K\subset G$ is compact and $U\subset G$ is open. With respect to the Chabauty topology, $\text{Sub}_{G}$ is a compact metrizable space (cf. Lemma E.1.1 of [3]). Moreover, a sequence $H_{n}\in\text{Sub}_{G}$ converges to $H\in\text{Sub}_{G}$ if and only if for any $h\in H$ there is a sequence $h_{n}\in H_{n}$ such that $h_{n}$ converges to $h$ and for any sequence $x_{k_{n}}\in H_{k_{n}}$, with $k_{n+1}>k_{n}$ and $x_{k_{n}}\to x$, we have $x\in H$ (cf. Lemma E.1.2 of [3]).
The group $G$ acts on SubG by conjugation: $(g,H)\mapsto gHg^{-1}$. An invariant random subgroup (hereafter IRS) is a probability measure $\mu$ on SubG that is invariant under the conjugation action of $G$ on SubG. We say that a Borel subset ${\mathcal{O}}$ of SubG supports an IRS $\mu$ if $\mu({{\mathcal{O}}})=1$: a Borel subset ${{\mathcal{O}}}$ supporting an IRS $\mu$ need not be the support of $\mu$ but $\overline{{\mathcal{O}}}$ contains the support of $\mu$. Dirac measure supported on a closed normal subgroup of a connected Lie group is an example of an IRS. Suppose $\Gamma$ is a lattice inside the connected Lie group $G$, i.e., $\Gamma$ is a discrete subgroup of $G$ such that $G/\Gamma$ carries a probability measure $\mu$ that is invariant under the natural action of $G$ on $G/\Gamma$. Then the map from $G$ to $\text{Sub}_{G}$ sending $g$ to $g\Gamma g^{-1}$ factors through $G/\Gamma$, and the push forward of $\mu$ by this map becomes a conjugation-invariant probability measure supported on conjugacy classes of $\Gamma$. So, invariant random subgroups can be thought of as a generalization of both normal subgroups and lattices.
Invariant random subgroups arise naturally as stabilizers of measure preserving actions of $G$ on a probability space $(X,\mu)$. By a result of Varadarajan ([14]), the stabilizer of almost every point of this action is closed in $G$. Moreover, the stabilizer map from $X$ to $\text{Sub}_{G}$ given by $x\to G_{x}$ is measurable. Then, one can push the measure $\mu$ forward to an IRS of $G$. In fact, every IRS can be obtained this way (see [1], [2] for more details).
In recent times, IRS has attracted many studies (see for examples [1] and [2]). In this article, we are interested in the existence of invariant random subgroups supported on certain orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. We start with the orbit of a semisimple Levi subgroup. Let $G$ be a connected Lie group with a closed semisimple Levi subgroup $L$; it may be noted that semisimple Levi subgroups of connected Lie groups need not be closed. But, Levi subgroups of linear groups are closed. Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$. Then $L_{G}=\{xLx^{-1}\mid x\in N\}$ where $N$ is the nilradical of $G$ (cf. Theorem 3.18.3 of [13]). An IRS $\mu$ is called a Levi IRS if $\mu(L_{G})=1$.
@@
In particular, the set of semisimple Levi subgroups $L_{G}$ is closed in SubG if and only if $L$ is normal.
@@
So, any connected Lie group $G$ admitting a Levi IRS is of the form $G=L\times S$ where $L$ is semisimple and $S$ is solvable. In particular, $G=\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS.
@@
Let $G$ be a connected real algebraic group and $B_{G}$ be the set of all Borel subgroups of $G$. Then $B_{G}$ is closed in SubG and there is an IRS supported on $B_{G}$ if and only if $G$ is amenable.
@@
Let $G=SO(3,\mathbb{R})\ltimes\mathbb{R}^{3}$. We identify $O(2,\mathbb{R})$ as a subgroup of $SO(3,\mathbb{R})$ by the map $\alpha\mapsto\begin{pmatrix}\alpha&0\cr 0&\det(\alpha)\end{pmatrix}$. Then $B=SO(2,\mathbb{R})\ltimes R^{3}$ is a Borel subgroup of $G$ and the normalizer of $B$ is $O(2,\mathbb{R})\ltimes\mathbb{R}^{3}$. Thus, $SO(3,\mathbb{R})$-invariant measure on the quotient space $SO(3,\mathbb{R})/O(2,\mathbb{R})$ induces an IRS on the set of Borel subgroups $B_{G}$ of $G$.
@@
Let $G$ be a connected real algebraic group and $D_{G}$ be the set of all maximal diagnolizable subgroups of $G$. Then there is an IRS supported on $D_{G}$ if and only if $D$ is central in $G$.
We have seen that the group $\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas, the group $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS. In particular, the orbit of a semisimple Levi subgroup is not closed in $\text{Sub}_{G}$ when $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. This leads to a natural question of studying the closure of the orbit under the conjugation action of a semisimple Levi subgroup. The same question can be asked for other orbits discussed above as well. In this article, we address these questions in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Note that, studying these questions is the same as investigating about conjugacy limits of various subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. The study of conjugacy limits is an interesting problem in its own right and has been an active area of research aimed at understanding the space $\text{Sub}_{G}$. For example, in [10], [11], the author has studied the conjugacy limits of diagonal subgroups in $\operatorname{SL}(3,\mathbb{R})$ and $\operatorname{SL}(n,\mathbb{R})$ respectively. One may also look at the references cited in these two papers for related works and their relation to geometry.
Given a closed subgroup $H$ of $G$, by a conjugacy limit of $H$ in $\text{Sub}_{G}$, we mean the limit of $(g_{n}Hg_{n}^{-1})$ for a sequence $(g_{n})$ in $G$. Note that $(g_{n}Hg_{n}^{-1})$ always has a convergent subsequence as $\text{Sub}_{G}$ is compact. The following theorem gives a complete description of all the limit points of the set $L_{G}$. We write any element of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ as $(g,v)$ with $g\in\operatorname{SL}(2,\mathbb{R})$ and $v\in\mathbb{R}^{2}$. Also, for any two elements $(g,v)$ and $(g^{\prime},v^{\prime})$, the group operation is defined as follows:
$(g,v)(g^{\prime},v^{\prime})=(gg^{\prime},v+g(v^{\prime})).$
We identify $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ as closed subgroups of $G$ by the maps $g\mapsto(g,(0,0))$ and $v\mapsto(I,v)$ where $I$ denotes the $2\times 2$ identity matrix. For any $c\in\mathbb{R}$, we define a unipotent subgroup $N^{+}_{c}$ by
$N^{+}_{c}=\left\{\big{(}\begin{pmatrix}1&x\\ 0&1\end{pmatrix},(cx,0)\big{)}:x\in\mathbb{R}\right\}.$
For $c=0$, we denote $N^{+}_{c}$ by $N^{+}$, which is identified (and isomorphic) to the group of unipotent upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$.
Theorem 1.7.
Any subgroup in $\overline{L_{G}}$ is either a Levi subgroup of $G$ or a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
Next, we consider the conjugacy limit of maximal compact subgroups, that is, the closure of $K_{G}=\{gKg^{-1}\mid g\in G\}$ for any maximal compact subgroup $K$ of $G=SL(2,\mathbb{R})\times\mathbb{R}^{2}$. For example, we can take $K=SO(2,\mathbb{R})$. For this purpose, we introduce the following subgroups of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$:
$V_{c}=\{(I,(t,ct))\mid t\in\mathbb{R}\},\ c\in\mathbb{R},~{}~{}V_{\infty}=\{(I,(0,t))\mid t\in\mathbb{R}\}$
and
$\tilde{N}^{+}=\left\{\pm\begin{pmatrix}1&s\\ 0&1\end{pmatrix}:s\in\mathbb{R}\right\}.$
It may be noted that $V_{c},c\in\mathbb{R}\cup\{\infty\}$ is the set of all one-dimensional subspaces of $\mathbb{R}^{2}$, the projective space ${\mathbb{P}}^{1}$.
Theorem 1.8.
The closure of $K_{G}$ consists of conjugates of the subgroups $V_{c},c\in\mathbb{R}\cup\{\infty\}$ and $\tilde{N^{+}}$, in addition to the conjugates of $K$.
The next result identifies limit points of the set of maximal diagonalizable subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Let $A=\left\{\begin{pmatrix}a&0\\ 0&\frac{1}{a}\end{pmatrix}\mid a\in\mathbb{R}^{+}\right\}$ and $D_{G}=\{gAg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}A$. We now introduce the following one-dimensional subgroups
$V_{(a,b,c)}=\left\{\left(\begin{pmatrix}1&at\\ 0&1\end{pmatrix},\left(ct+\frac{abt^{2}}{2},bt\right)\right)\mid t\in\mathbb{R}\right\}$
for any $a,b,c\in\mathbb{R}$. It may be noted that $V_{a,b,c}$’s are the one-parameter subgroups in the $3$-dimensional Heisenberg group and $V_{(1,0,c)}=N^{+}_{c}$, $V_{(0,1,c)}=V_{c}$ and $V_{(0,0,1)}=V_{\infty}$.
Theorem 1.9.
The closure of $D_{G}$ consists of conjugates of $V_{(r_{1},r_{2},r_{3})}$ for any $r_{1},r_{2},r_{3}\in\mathbb{R}$ with $r_{1}=1$ or $r_{3}=1$ or $r_{1}=0$ and $r_{2}=1$ in addition to the conjugates of the diagonal group $A$.
Let $B$ denote the group of all upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$. Then $B$ is a Borel subgroup of $\operatorname{SL}(2,\mathbb{R})$ and $B\ltimes\mathbb{R}^{2}$ is a Borel subgroup of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Then, using the Iwasawa decomposition in $\operatorname{SL}(2,\mathbb{R})$, it is easy to see that, for conjugacy limits of $B\ltimes\mathbb{R}^{2}$, we only need to consider conjugation by elements of $K=\operatorname{SO}(2,\mathbb{R})$, hence conjugates of $B$ form a closed subset (but by Theorem 1.4 no IRS is supported on the conjugates of $B$ as $G$ is not amenable). We therefore consider conjugacy limits of $B$ inside $G$ and identify all possible conjugacy limits of $B$
Theorem 1.10.
The closure of $\{gBg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $B$, $N^{+}\ltimes\{(t,0)\mid t\in\mathbb{R}\}~{}(\simeq\mathbb{R}^{2})$ and $\mathbb{R}^{2}$.
The next result is concerned about the conjugacy limits of the unipotent subgroup $N^{+}$.
Theorem 1.11.
The closure of $\{gN^{+}g^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $N^{+}$ and $V_{0}$.
@@ 2. Preliminary
In this section, we prove certain technical results which are needed to discuss the existence or non-existence of IRS on orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. Given a finite-dimensional vector space $V$, let ${\rm Gr}_{k}(V)$ be the be the Grassnannian manifold of $k$-dimensional subspaces of $V$. Equip $V$ with an inner product, and for a subspace $W$ of $V$, let $P_{W}$ denote the orthogonal projection onto the subspace $W$. We define a metric on ${\rm Gr}_{k}(V)$ using the operator norm ||.||||.||| | . | | of $P_{W}$’s as follows: for $W,W^{\prime}\in{\rm Gr}_{k}(V)$,
$d(W,W^{\prime})=||P_{W}-P_{W^{\prime}}||.$
The following Lemma is a comparison between the Grassmannian topology and the Chabauty topology of ${\rm Gr}_{k}(V)$ seen as a subset of $\text{Sub}_{V}$.
@@
Let $V$ be a finite-dimensional vector space and ${\rm Gr}_{k}(V)$ be the Grassmannian manifold of $k$-dimensional subspaces of $V$. Let $\mathfrak{i}\colon{\rm Gr}_{k}(V)\to{\rm Sub}_{V}$ be the canonical inclusion. Then ${\rm Gr}_{k}(V)$ is homeomorphic to its image. In particular, the set of all subspaces of dimension $k$ is a closed set in Chaubauty topology.
@@
Suppose $W_{n}\to W$ in ${\rm Gr}_{k}(V)$. This implies that $||P_{W_{n}}-P_{W}||\rightarrow 0$ as $n\rightarrow\infty$. Now let $w\in W$, we want to show that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$. Let $w_{n}=P_{W_{n}}(w)$. Then clearly $w_{n}\in W_{n}$ and
$||w_{n}-w||=||P_{W_{n}}(w)-P_{W}(w)||\leq||P_{W_{n}}-P_{W}||||w||\rightarrow 0$
as $n\rightarrow\infty$. On the other hand assume that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$, we want to show that $w\in W$. It is enough to show that $P_{W}(w)=w$. Since $w_{n}\rightarrow w$ and $P_{W_{n}}\rightarrow P_{W}$, it follows that $P_{W_{n}}(w_{n})\rightarrow P_{W}(w)$. But as $P_{W_{n}}(w_{n})=w_{n}$, and $w_{n}\rightarrow w$, it follows that $P_{W}(w)=w$. Thus, $W_{n}\to W$ in SubV. This proves that the map $\pi$ is continuous. Since ${\rm Gr}_{k}(V)$ is compact, ${\rm Gr}_{k}(V)$ is homeomorphic to $\mathfrak{i}({\rm Gr}_{k}(V))$. ∎
It is known that if a sequence of closed subgroups $H_{n}$ of a Lie group $G$ converges to a closed subgroup $H$, then dim$(H)\geq{\rm limsup}~{}\text{dim}(H_{n})$ (cf. Theorem 1 of [15]). However, in the case of closed subgroups of thev same dimension and in the case of exponential Lie groups, we have the following.
@@
Let $G$ be a connected Lie group and $H_{n}\to H$ in SubG. Let $\mathcal{G}$, $\mathcal{H}_{n}$ and $\mathcal{H}$ be Lie algebras of $G$, $H_{n}$ and $H$ respectively.
  1. (1)
    If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
  2. (2)
    If $G$ is an exponential Lie group and $H_{n}$’s are connected, then $H$ is also connected, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG and ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
@@
  2. (2)
    It may be noted that $H$ need not always be connected even when $H_{n}$ are all connected and have the same dimension: e.g., cases $1$ and $2$ of Section $5.3$ provide a sequence of conjugate subgroups whose limit point has two connected components.
@@
Let $V$ be a limit of $\mathcal{H}_{n}$ in SubG. By passing to a subsequence, we may assume that $\mathcal{H}_{n}\to V$ in SubG. Then $V$ is a vector subspace. In is easy to see that it is a Lie subalgebra of $\mathcal{G}$ as well. For $v\in V$, there are $v_{n}\in\mathcal{H}_{n}$ such that $v_{n}\to v$, hence $tv_{n}\to tv$ for all $t\in\mathbb{R}$. Let $\exp\colon\mathcal{G}\to G$ be the exponential map. Then $\exp(tv_{n})\to\exp(tv)$ for all $t\in\mathbb{R}$. Since $H_{n}$ converges to $H$, $\exp(tv)\in H$ for all $t\in\mathbb{R}$. This implies that $v\in\mathcal{H}$. Thus, $V\subset\mathcal{H}$.
Assume that $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$. Then $\text{dim}(\mathcal{H}_{n})=\text{dim}(\mathcal{H})$ for all $n$. By Lemma 2.1, $\text{dim}(V)=\text{dim}(\mathcal{H})$. Since $V\subset\mathcal{H}$, we get that $V=\mathcal{H}$. Thus, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
Assume that $\exp\colon\mathcal{G}\to G$ is a diffeomorphism and $H_{n}$ are connected. Then $\exp\colon\mathcal{H}_{n}\to H$ is also a diffeomorphism for all $n$. Let $h\in H$. Since $H_{n}\to H$, there are $h_{n}\in H_{n}$ such that $h_{n}\to h$. Let $X_{n}\in\mathcal{H}_{n}$ and $X\in\mathcal{H}$ be such that $\exp(X_{n})=h_{n}$ and $\exp(X)=h$. Since $\exp$ is a diffeomorphism, we get that $X_{n}\to X$. Since $\mathcal{H}_{n}\to V\subset\mathcal{H}$, we get that $X\in\mathcal{H}$. Thus, $\exp(\mathcal{H})=H$. Therefore, $H$ is connected. Since $\exp$ is a diffeomorphism, we also get that $\mathcal{H}_{n}\to\mathcal{H}$ and hence $V=\mathcal{H}$. Thus, ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
The result of the following kind is known as Furstenberg’s Lemma (cf. Section 5, [7]) and it has been proved for minimally almost periodic groups in Lemma 3 of [6] and when $\Gamma$ (as in Lemma 2.4) is certain semisimple analytic group in Proposition 5.1 of [7]. Here we provide the following using a result of [5].

### v2

# Invariant random subgroups on certain orbits

@@ Abstract.
Let $G$ be a connected Lie group and $\text{Sub}_{G}$ be the space of closed subgroups of $G$ equipped with the Chabauty topology. In this article, we investigate the existence of invariant random subgroups of $G$ supported on various orbits of the conjugation action of $G$ on $\text{Sub}_{G}$.
@@ 1. Introduction
Let $G$ be a connected (real) Lie group and $\text{Sub}_{G}$ be the space of (closed) subgroups of $G$ equipped with the Chabauty topology introduced by Claude Chabauty in [6]. The Chabauty topology on $\text{Sub}_{G}$ is generated by the subsets $\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$, $\{H\in{\rm Sub}_{G}\mid H\cap U\not$$\emptyset\}$, where $K\subset G$ is compact and $U\subset G$ is open. With respect to the Chabauty topology, $\text{Sub}_{G}$ is a compact metrizable space (cf. Lemma E.1.1 of [4]). Moreover, a sequence $H_{n}\in\text{Sub}_{G}$ converges to $H\in\text{Sub}_{G}$ if and only if for any $h\in H$ there is a sequence $h_{n}\in H_{n}$ such that $h_{n}$ converges to $h$, and for any sequence $x_{k_{n}}\in H_{k_{n}}$, with $k_{n+1}>k_{n}$ and $x_{k_{n}}\to x$, we have $x\in H$ (cf. Lemma E.1.2 of [4]).
The group $G$ acts on $\text{Sub}_{G}$ by conjugation: $(g,H)\mapsto gHg^{-1}$. An invariant random subgroup (hereafter IRS) is a probability measure $\mu$ on $\text{Sub}_{G}$ that is invariant under the conjugation action of $G$ on $\text{Sub}_{G}$. Dirac measure supported on a closed normal subgroup is an IRS. Also, if $\Gamma$ is a lattice in $G$, i.e., $\Gamma$ is a discrete subgroup of $G$ such that $G/\Gamma$ carries a $G$-invariant probability measure $\mu$, then the push forward of $\mu$ under the map $g$ to $g\Gamma g^{-1}$ defines an IRS on $G$. Thus, invariant random subgroups can be thought of as a generalization of both normal subgroups and lattices. The study of invariant random subgroups and ergodic theory are deeply intertwined as an IRS can be realized as stabilizers of a measure-preserving action of the group (see [1] and [3] for more details).
We say that a Borel subset $\mathcal{O}$} of $\text{Sub}_{G}$ supports an IRS $\mu$ if $\mu({{\mathcal{O}}})=1$: a Borel subset $\mathcal{O}$}} supporting an IRS $\mu$ need not be the support of $\mu$ but $\overline{{\mathcal{O}}}$ contains the support of $\mu$.
In recent times, IRS has attracted many studies (cf. [1], [2], [3], [5], [9] and the references therein). In this article, we are interested in the existence of invariant random subgroups supported on certain orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. We first consider the orbit of semisimple Levi subgroups. Let $G$ be a connected Lie group with a closed semisimple Levi subgroup $L$; it may be noted that semisimple Levi subgroups of connected Lie groups need not be closed, but, Levi subgroups of linear groups are closed. Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$. Then $L_{G}=\{xLx^{-1}\mid x\in N\}$ where $N$ is the nilradical of $G$ (cf. Theorem 3.18.3 of [13]). An IRS $\mu$ is called a Levi IRS if $\mu(L_{G})=1$.
@@
In particular, the set of semisimple Levi subgroups $L_{G}$ is closed in $\text{Sub}_{G}$ if and only if $L$ is normal.
@@
Thus, any connected Lie group $G$ admitting a Levi IRS is of the form $G=L\times S$ where $L$ is semisimple and $S$ is solvable. In particular, for $n\geq 2$, $\operatorname{SL}(n,\mathbb{R})\times\mathbb{R}^{n}$ admits a Levi IRS, whereas $\operatorname{SL}(n,\mathbb{R})\ltimes\mathbb{R}^{n}$ does not admit a Levi IRS.
@@
Let $G$ be a connected real algebraic group and $B_{G}$ be the set of all Borel subgroups of $G$. Then $B_{G}$ is closed in $\text{Sub}_{G}$, and there is an IRS supported on $B_{G}$ if and only if $G$ is amenable.
@@
Let $G=SO(3,\mathbb{R})\ltimes\mathbb{R}^{3}$. We identify $O(2,\mathbb{R})$ as a subgroup of $SO(3,\mathbb{R})$ by the map $\alpha\mapsto\begin{pmatrix}\alpha&0\cr 0&\det(\alpha)\end{pmatrix}$. Then $B=SO(2,\mathbb{R})\ltimes R^{3}$ is a Borel subgroup of $G$ and the normalizer of $B$ is $O(2,\mathbb{R})\ltimes\mathbb{R}^{3}$. Thus, any $SO(3,\mathbb{R})$-invariant measure on the quotient space $SO(3,\mathbb{R})/O(2,\mathbb{R})$ induces an IRS on the set of Borel subgroups $B_{G}$ of $G$.
@@
Let $G$ be a connected real algebraic group and $D_{G}$ be the set of all maximal diagnolizable subgroups of $G$. Then the following are equivalent:
  1. (1)
    there is an IRS supported on $D_{G}$;
  2. (2)
    $D$ is central in $G$;
  3. (3)
    $D$ is normal in $G$.
Moreover, if (1) holds, then any Levi subgroup of $G$ is compact and the solvable radical is a compact extension of its nilradical.
@@ 2. Preliminary
In this section, we now collect a few results that connect the dynamics of $G$ on $\text{Sub}_{G}$ with projective linear actions on Grassmannians. Given a finite-dimensional vector space $V$, let ${\rm Gr}_{k}(V)$ be the Grassnannian manifold of $k$-dimensional subspaces of $V$.
We include a proof of the following lemma because we are unable to locate it in the following form that compares the Grassmannian topology and the Chabauty topology of ${\rm Gr}_{k}(V)$ seen as a subset of $\text{Sub}_{V}$. Recall that the metric on ${\rm Gr}_{k}(V)$ is given by
[…]
for $W,W^{\prime}\in{\rm Gr}_{k}(V)$, with $P_{W}$ and $P_{W}^{\prime}$ being the orthogonal projections onto the subspaces $W$ and $W^{\prime}$ respectively. One may look at [9] for the measurability of the map.
@@
Let $V$ be a finite-dimensional vector space and ${\rm Gr}_{k}(V)$ be the Grassmannian manifold of $k$-dimensional subspaces of $V$. Let $\mathfrak{i}\colon{\rm Gr}_{k}(V)\to{\rm Sub}_{V}$ be the canonical inclusion. Then ${\rm Gr}_{k}(V)$ is homeomorphic to its image. In particular, the set of all subspaces of dimension $k$ is a closed set in the Chaubauty topology.
@@
Suppose $W_{n}\to W$ in ${\rm Gr}_{k}(V)$. This implies that $P_{W_{n}}-P_{W}||\rightarrow 0$ as $n\rightarrow\infty$. Now let $w\in W$, we want to show that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$. Let $w_{n}=P_{W_{n}}(w)$. Then clearly $w_{n}\in W_{n}$ and
[…]
as $n\rightarrow\infty$. On the other hand assume that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$, we want to show that $w\in W$. It is enough to show that $P_{W}(w)=w$. Since $w_{n}\rightarrow w$ and $P_{W_{n}}\rightarrow P_{W}$, it follows that $P_{W_{n}}(w_{n})\rightarrow P_{W}(w)$. But as $P_{W_{n}}(w_{n})=w_{n}$, and $w_{n}\rightarrow w$, it follows that $P_{W}(w)=w$. Thus, $W_{n}\to W$ in SubV. This proves that the map $\mathfrak{i}$ is continuous. Since ${\rm Gr}_{k}(V)$ is compact, ${\rm Gr}_{k}(V)$ is homeomorphic to $\mathfrak{i}({\rm Gr}_{k}(V))$. ∎
[…]
@@
Let $G$ be a connected Lie group and $H_{n}\to H$ in $\text{Sub}_{G}$. Let $\mathcal{G}$, $\mathcal{H}_{n}$ and $\mathcal{H}$ be Lie algebras of $G$, $H_{n}$ and $H$ respectively. If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
@@
Let $V$ be a limit of $\mathcal{H}_{n}$ in SubG. By passing to a subsequence, we may assume that $\mathcal{H}_{n}\to V$ in SubG. Then $V$ is a vector subspace, and it is easy to see that $V$ is a Lie subalgebra of $\mathcal{G}$ as well. For $v\in V$, there are $v_{n}\in\mathcal{H}_{n}$ such that $v_{n}\to v$, hence $tv_{n}\to tv$ for all $t\in\mathbb{R}$. Let $\exp\colon\mathcal{G}\to G$ be the exponential map. Then $\exp(tv_{n})\to\exp(tv)$ for all $t\in\mathbb{R}$. Since $H_{n}$ converges to $H$, $\exp(tv)\in H$ for all $t\in\mathbb{R}$. This implies that $v\in\mathcal{H}$. Thus, $V\subset\mathcal{H}$.
If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\text{dim}(\mathcal{H}_{n})=\text{dim}(\mathcal{H})$ for all $n$. By Lemma 2.1, $\text{dim}(V)=\text{dim}(\mathcal{H})$. Since $V\subset\mathcal{H}$, we get that $V=\mathcal{H}$. Thus, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
3. On the support of random subgroups invariant under unipotent transformations
In this section we consider invariant measures on Grassmannian for projective linear actions and first prove a result on the support of the invariant measure. The result of the following kind is known as Furstenberg’s Lemma (cf. Section 5, [9]) and it has been proved for minimally almost periodic groups in Lemma 3 of [8] and when $\Gamma$ (as in Lemma 3.1) is certain semisimple analytic group in Proposition 5.1 of [9]. Here we provide the following using a result of [7].